OpenStudy (crashonce):

If f(x) = f(-x), then f(x) is called an even function. Which of the following are even functions? 1)f(x) =x^-2 2)f(x)=x^4 3)f(x)=2x^3

OpenStudy (crashonce):

@ganeshie8 @mathmath333

OpenStudy (jhannybean):

plug in values.

OpenStudy (jhannybean):

example: \(f(2) = (2)^2 = ~?\), \(f(-2) = (-2)^22 = ~?\)

OpenStudy (jhannybean):

ahaha typo.

OpenStudy (jhannybean):

\[f(-2) = (-2)^2 = ~?\]

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{f(x)=x^{-2}-----\color{red}{(1)}(given)\\~\\ f(-x)=(-x)^{-2}\\~\\ f(-x)=\dfrac{1}{(-x)^{2}}\\~\\ f(-x)=\dfrac{1}{(x)^{2}}\\~\\ f(-x)=(x)^{-2}-----\color{red}{(2)}\\~\\~\\ f(x)=f(-x) }\end{align}\)

OpenStudy (zzr0ck3r):

\(f(x) = x^{-2}\) we have \(f(-x)=\frac{1}{(-x)^2}=\frac{1}{x^2}=f(x)\)

OpenStudy (crashonce):

so is it just the first one

OpenStudy (zzr0ck3r):

no

OpenStudy (jhannybean):

No guessing >:o

OpenStudy (crashonce):

I don't quite get how it wokrs

OpenStudy (crashonce):

@Jhannybean

OpenStudy (mathmath333):

u have to substitute \(-x\) for \(x\) and see if \(f(x)=f(-x)\)

OpenStudy (jhannybean):

^^^^

OpenStudy (mathmath333):

refer this http://www.mathsisfun.com/algebra/functions-odd-even.html

OpenStudy (jhannybean):

"x" is just a value, a number, all you have to do is plug in a number.

OpenStudy (crashonce):

I still don't know how it works, how should I approach this

OpenStudy (jhannybean):

So in functions that have an odd exponent, if you plug in a negative value for x, you will get a negative result. Example : \((-1)^{3} = (-1)(-1)(-1) = -1\)

OpenStudy (crashonce):

so can u take me through one answer option by showing me if it is or isn't

OpenStudy (unklerhaukus):

If the polynomial has only even powers (+/-), it is an even function, if the polynomial only has odd powers (+/-), it is an odd function.

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{f(x)=x^{4}-----\color{red}{(1)}(given)\\~\\ f(-x)=(-x)^{4}\\~\\ f(-x)=(-1)^4(x)^{4}\\~\\ f(-x)=(1)x^{4}\\~\\ f(-x)=x^{4}-----\color{red}{(2)}\\~\\ \color{red}{(1)=(2)}\\~\\ f(x)=f(-x)\\ even }\end{align}\)

OpenStudy (crashonce):

ok ive worked out the third is odd am I right?

OpenStudy (mathmath333):

\[\Huge \checkmark\]